OpenStudy (anonymous):

Find C and a so that f(x)=Ca^x satisfies the given conditions. f(0)=17, f(1)=51

OpenStudy (asnaseer):

ok, so you are given:\[f(x)=Ca^x\](1) f(0)=17, this means when x=0, f(x)=17, so we get:\[17=Ca^0=C*1=C\]so now we know C=17 (2) f(1)=51, this means when x=1, f(x)=51, so we get:\[51=Ca^1=Ca=17a\]\[a=\frac{51}{17}=3\]